OpenStudy (anonymous):

Solve 5x + 14 - 2x = 9 - (4x + 2) 1) x =3/7 2) x = - 1 3) x = 6 4) x = 8

jhonyy9 (jhonyy9):

5x+14-2x=9-(4x+2) --- 3x+14=9-4x-2 --- 7x=-7 --- x= -1 2). x= -1